A force of 9.0N is applied to a body at rest on a smooth horizontal surface. If the body has a mass of 18.0kg and it moves in the direction of the force,find the work done in the first four seconds and the fourth second
1
Expert's answer
2019-03-18T12:41:33-0400

The body acceleration a = F/m, then its velocity v = at = Ft/m, and kinetic energy K = mv2/2 = (Ft)2/2m. The work done A equals to K, so A1 = (9*4)2/2*18 = 36 (J), A2 = (9*4)2/2*18 - (9*3)2/2*18 = 15.8 (J).


   Finally, A1 = 36 J, A2 = 15.8 J.